Question 5 3 pts Consider two assets, A and B, with the following equally likely rates of return: A will return either 3%, 6%, or 3.5%. B will return either 2.5%, 8%, or 7.5 Asset A Asset B 3.5% 7.5% 3.0% 6.0% 8.0% 2.5% Find the expected rates of return for assets A and B. Which asset would an investor, who cares only about expected return and risk and can choose only one or the other, prefer? A is better Not Clear Equal B is better O
Answer #1

A is better

Expected rate of return for A=1/3*(3%+6%+3.5%)=4.167%

Expected rate of return for B=1/3*(2.5%+8%+7.5%)=6%

Standard Deviation of A=sqrt(1/3*(3%-4.167%)^2+1/3*(6%-4.167%)^2+1/3*(3.5%-4.167%)^2)=1.31%

Standard Deviation of B=sqrt(1/3*(2.5%-6%)^2+1/3*(8%-6%)^2+1/3*(7.5%-6%)^2)=2.48%

Return to Risk of A=4.167%/1.31%=3.1809

Return to Risk of B=6%/2.48%=2.4194

Hence choose A